Use the H-H equation and Table 2.6 to do the following calculations. For the solutions with...

  1. Use the H-H equation and Table 2.6 to do the following calculations. For the solutions with the following pH values, calculate the ratio of the concentrations of the base and the acid forms of the substances listed.  
    1. for an acetic acid/acetate solution with a pH of 2.66:   [A-]/[HA] =  ______________
    2. for a phenol/phenolate solution with a pH of 10.24:  [A-]/[HA] =  _______________
    3. for a phosphate buffered solution with a pH of 6.55:  [A-]/[HA] =   _______________
Answer #1

According to H-H equation:

-pKa = -pH + log([A-]/[HA])

a. pH = 2.66, pKa of acetate solution: 4.76

[A-]/[HA] = 10^(-4,76+2.66) = 0.0079

b. pH = 10.24, pKa of phenol: 10

[A-]/[HA] = 10^(-10+10.24) = 1.74

c. pH = 6.55, pKa of phosphate buffered solution: 6.85

[A-]/[HA] = 10^(-6.85+6.55) = 0.501